OpenStudy (anonymous):

log(⁷√10000)= 4/7 ????

OpenStudy (anonymous):

log (10000)^1/7 now breakdown 10000=10x10x10x10 =10^4 so log(10^4)^1/7 which gives you log(10^4/7) the log10=1 in this case it is 4/7

OpenStudy (anonymous):

thanks